At the close of 2017, the financial statements of CLEO Corp. were as follows. All figures in the financial statements are in thousands ($000) CLEO Corp. Balance Sheet For the vear ended 12/31/2017 12/31/2016 12/31/2017 ASSETS Cash 500 $ 200 CLEO Corp Income Statement For the year ending 12/31/2017 Accounts Receivable Inventory CURRENT ASSETS 6.250 5.180 $11.930 7.300 6.470 $13.970 Sales COGS Gross Margin $22.560 11.506 $11.054 Fixed Assets Expense $5.332 Gross Accum. Deprec. Net $7.500 (2.400 $5,100 $9.000 (3.100 $5.900 Depreciation EBIT Interest EBT Tax Net Income 700 $5.022 1.180 $3.842 1.537 $2.305 TOTAL ASSETS $17.030 $19.870 LIABILITIES Accounts Payable Accruals CURRENT LIABILITIES Long-term Debt Equity TOTAL CAPITAL TOTAL LIABILITIES AND EQUITY $1,860 850 $2.710 $11,320 3.000 $14.320 $2.210 $2.430 $12.335 5.105 $17.440 $17,030 $19.870 In addition, CLEO Corp. paid dividends of $1.2M and sold new stock valued at $1.0M during 2017. Required: Prepare the SCF. (Show all your work!)
Answer #1

Answer:

CLEO Corp Statement of Cash Flows For the Year ended December 31, 2017 Operating Activities Net Income Adiustment to convert net income to a cash basis Add, Depreciation Add, Decrease/(Increase) in Invento Add, Decrease/(Increase) in Account Receivables Add, Increase /(Decrease) in Account payable Add, Increase/(Decrease) in Accrued Liabilities Cash flow from Operating Activities (in $ 000 $2,305 $700 ($1,290) (S1,050) $350 630)($1,920) $385 Investing Activities Purchase of Fixed assets Cash flow from Investing Activities ($1,500) $1,500) Financing Activities Long Term borrowings Proceeds from Issues of Shares Dividend Paid Cash flow from Financing Activities $1,015 $1,000 ($1,200) $815 S300) $500 Change in Cash flow Beginning Cash Balance Ending Cash Balance $200

Workings:

Increase in inventory = 6470 - 5180 = 1290

Increase in account receivable = 7300 - 6250 = 1050

Increase in account payable = 2210 - 1860 =350

Decrease in accrued liabilities = 850 - 220 = 630

Purchase of fixed assets = Gross value of assets 2017 - Gross value of assets 2016 =9000 -7500 = 1500

Long term borrowing = Long term borrowing 2017 - Long term borrowing 2016 = 12335 - 11320 =1015